if thirty pounds mixed nuts sells for $6.87 per pound. the mixture is obtained from two kinds of nuts, peanuts priced at $5.70 per pound and cashews at $8.70 per pound.how many pound of each variety of nuts are used in the mixture?
Answer by mananth(16949)

if thirty pounds mixed nuts sells for $6.87 per pound. the mixture is obtained from two kinds of nuts, peanuts priced at $5.70 per pound and cashews at $8.70 per pound.how many pound of each variety of nuts are used in the mixture?
..
peanuts priced at $5.70 per pound: let the quantity be x lbs
cashews at $8.70 per pound: this will be 30-x lbs
Mixture $6.87 per pound. 30 lbs
...
5.70x+8.70(30-x)=30*6.87
5.70x+261-8.70x =206.10
-3x=-54.9
/-3
x=18.3 lbs peanuts
11.7 lbs = cashews
m.ananth@hotmail.ca